40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5141     Entry time: Mon Aug 8 15:20:49 2011
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: inspection on ETMY (round 1) 

Since ETMY have been showing some strange behaviors we deeply inspect the ETMY suspension.

Here is a brief review of the ETMY suspension and a brief status update of the inspection so far.

The inspection is still ongoing.

 

(Review : How wrong ?)

First of all, let us summarize what were the observed phenomena on the ETMY suspension :

   1. unknown low frequency noise covering frequency range from 0.1 to 3 Hz in all of four face sensors (#5025, #5029).

   2. LR sensor showed a very broad bounce peak at ~ 17 Hz (#5025).

   3. The sign of some of the sensors are flipped.

   4. The control gains had to be higher than those of ETMX by 10-100 (#5025).

 

(Status update : Noise spectra)

Currently the ETMY suspension is sitting on the north side of the table for the inspection.

We took dark noise of the OSEMs when the OSEMs were taken off from the tower and put on the table.

The plot below is an example of the LR sensor spectra. Note that the whitening filters have been always ON.

Screen_shot_2011-08-08_at_15.08.23.png

The black curve is the dark noise when the sensor was off from the tower.

The blue curve is the free swinging spectrum newly taken today.

The red curve is the free swinging spectrum (damped spectrum ?) on 25th of July, this was still in vacuum.

 

The dark noise is below the free swinging spectrum from 0.2 - 30 Hz, which looks reasonable

The most interesting thing is that the free swinging spectrum became better in low frequency (below 3 Hz)

from the one measured in vacuum.

It needs more investigation to answer the reason why it happened.

Note that before we moved the tower to the current position, we looked at the OSEM-magnets relations, and found nothing was touching.

 

ELOG V3.1.3-